Shift summary for 08/31/99 Owl shift

Juliana Whitmore (jaws@kpasa)
Tue, 31 Aug 1999 06:57:44 -0500

Shift Summary for Owl shift, 31/Aug/99

On Shift: Carlos, Julie

Hours of beam available at all intensities: 7 hrs
Hours of beam available at nominal intensity: 7 hrs
Hours of E832 data (special runs not included): 6.4 hrs
Hours of special runs (laser scans, muon runs...): 0.1 hr

Accounting of lost time: Tape change; NM4AN polarity flip (NEG to POS);
KFRENDing.

========================================================================
Runs ended this shift:
Run 14379 - 24.0M events - BEAM_832
Run 14380 - 125k events - BEAM_MUON

Tapes brought to FCC this shift:
(including tapes for online split)

QK*357-362, QKB124-QKB125

========================================================================
Accesses during shift:None

========================================================================
Overview of shift:
Smooth E832 datataking. Muon run and NM4AN magnet polarity flip.
Labelled tapes.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14379 00:00 Inherit smooth E832 data taking.

00:34 Run ends because tapes fill up. Zero NM4AN (NEGtoPOS).
Load new tapes (QK*363-5, QKB126).

00:40 Asked to install new ADGN constants. Done.

14380 00:47 Muon run. KTEVANA errors on 1st spill. KFREND. End
after a few spills.

00:50 NM4AN magnet polarity flip (AN832UP). Now positive.
Notified to update CSPD. Pedestal histogram looks
fine, so we decide not to update now.

14381 01:02 Start BEAM832 run. KTEVANA errors on first spill.
KFREND.

21 01:32 KTEVANA errors - missing FFFF word stream 5. KFREND.

31 01:45 KTEVANA errors - missing FFFF word stream 5. KFREND.

138-142 04:10 KTEVANA errors - missing FFFF word stream 5. KFREND.
Lose permit for stream 5. KFREND. Ordinal # mismatch
for CsI. Reset pipeline. KFREND.

05:40 Onsplit requesting tape mount - QKA022 loaded.

229 06:10 Val alarm. Laser RMS too wide - channel 1279. Goes
away after 2 spills.

Phoned MCR to complain that the intensity has been
slowly dropping off over the past hour. Max says he'll
look into it. Intensity is back up w/i a spill or two.

06:15 Ed phones. He's on his way out. He wants to remake
level 3 before the next run (at tape change).

237 06:23 KTEVANA errors - unknown crate stream 5. KFREND.

249 06:40 SEM too high - 7.2E12. No deadtime alarms, so we don't
phone MCR to complain.

06:50 Max calls to say that he's played with the splits and
is able to get us more beam but that we should expect
more pulse to pulse variation (6.8-7.2E12). Things
look ok (no deadtime), so we'll take what we can get.

07:00 Hand over to day crew.

========================================================================
Follow up Notes on Previous Shift

o When tapes are full, do muon run and set NM4AN polarity to positive. DONE.

========================================================================
Items to Follow up on Next Shift

o if we lose beam for >= 40 minutes, and the DA works, do a laser scan.

o keep your eyes on channel 450 (see log book p. 103). If it remains stable,
and we do a laser calibration, maybe we don't need to swap the divider
(ask Ed). If it is unstable, page Julie and Ed.

o Ed wants to remake level 3 before the next run (at tape change).

o trg#summary shows RC10 trigger source is bad. RSK says not to worry
since the extra triggers sources are o.k. Elizabeth has been notified.